OpenStudy (cherie_magee):

if a line contains the point (0,-1) and has a slope of 2, then what point lies on the line?

OpenStudy (nurali):

we first determine the equation of the line using the slope-intercept form y = mx + b where m = slope b = y-intercept given a slope of 2, the equation of the line would initially look like this: y = 2 x + b to compute for b, we substitute the values of point ( 0, -1 ) to the equation -1 = ( 2 ) ( 0 ) + b -1 = 0 + b -1 = b b = -1 the equation of the line is therefore y = 2 x - 1 to determine which of the following points also lies on the line, simply substitute the value of x and see if the resulting y is the same as the y of the point in question. a) ( 2, 1 ) y = ( 2 ) ( 2 ) - 1 = 4 - 1 = 3 does not lie on the line b) ( 1, 1 ) y = ( 2 ) ( 1 ) - 1 = 2 - 1 = 1 this point lies on the line c) ( 0, 1 ) y = ( 2 ) ( 0 ) - 1 = 0 - 1 = -1 does not lie on the line